OpenStudy (anonymous):

Hey, how can I simplify this? ((sqrt(x))-(1/(sqrt(x)))/(1-(1/(sqrt(x))))

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

\[(\sqrt{x}-1/\sqrt{x})/(1-1\sqrt{x)}\]

OpenStudy (dumbcow):

\[\frac{\sqrt{x}-\frac{1}{\sqrt{x}}}{1-\frac{1}{\sqrt{x}}}\] ?

OpenStudy (anonymous):

yes, I need to simplify that to \[1+\sqrt{x}\]

OpenStudy (anonymous):

I just don't know how :(

OpenStudy (dumbcow):

combine fractions on top first \[\sqrt{x} - \frac{1}{\sqrt{x}} = \frac{x-1}{\sqrt{x}}\]

OpenStudy (dumbcow):

then combine fractions on bottom \[1-\frac{1}{\sqrt{x}} = \frac{\sqrt{x}-1}{\sqrt{x}}\]

OpenStudy (dumbcow):

use rule when dividing fractions Flip and Multiply \[=\frac{x-1}{\sqrt{x}}*\frac{\sqrt{x}}{\sqrt{x}-1}\]

OpenStudy (dumbcow):

sqrt(x) cancel remember \[x^{2} -1 = (x-1)(x+1)\] so \[x-1 = (\sqrt{x}-1)(\sqrt{x}+1)\]

OpenStudy (anonymous):

Wow Thank you so much.

OpenStudy (dumbcow):

your welcome

OpenStudy (anonymous):

How do I place this as a "good answer"?

OpenStudy (dumbcow):

click on blue button to right of your name

OpenStudy (anonymous):

?

OpenStudy (dumbcow):

where it says "0medals for this answerer"

OpenStudy (dumbcow):

maybe it doesnt show up for you yet it this first time on

OpenStudy (anonymous):

Yeah, maybe.

OpenStudy (dumbcow):

heres one for you though

OpenStudy (anonymous):

Thank you!